# A New Hope: Contextual Privacy Policies for Mobile Applications and An Approach Toward Automated Generation

### 22 Feb 2024 v1

@@ 1 Introduction
Privacy policies have become the most prevalent privacy notice approach for mobile applications, with the aim of protecting users’ privacy and digital security [54, 39, 88, 56, 65, 29, 57, 37, 71]. On the contrary, application users often find themselves in a privacy paradox, specifically, Digital Resignation. This term refers to a state where users desire to understand and control the information that digital entities own about them, but feel powerless to do so [50]. A study has found that the average length of privacy policies is around 4,000 words for 75 prominent mobile applications and websites, requiring an estimated 16 minutes to read for an average adult [34]. Consequently, according to a survey conducted by Obar and Oeldorf-Hirsch [83], a significant portion of users, approximately 74%, opt for the “quick join” clickwrap option without engaging with privacy policies, primarily due to the lengthy and detailed nature of these documents. These numbers indicate that privacy policies have become increasingly difficult to read and understand in a timely manner, exacerbating the prevalence of digital resignation among users.
@@
In summary, our proposed framework extends the automated generation of CPPs to mobile applications, which are the digital platforms users spend the majority of time on [36, 67, 33]. The key contributions are:
@@ 2 Background
Mobile applications provide users with unparalleled convenience, multiple entertainment options, and various services. At the same time, they are responsible for the collection and exchange of substantial amounts of personal information, raising concerns surrounding data privacy. In response to the potential risks of privacy violations, the U.S. Department of Health, Education, and Welfare proposed a set of Fair Information Practice Principles (FIPP) in 1973 [7, 68]. Later, the Federal Trade Commission adopted these principles as “the notice and choice privacy framework” [83], which is commonly regarded as the foundation for modern privacy regulations such as the European General Data Protection Regulation (GDPR) [8], the California Consumer Privacy Act (CCPA) [4], and the Australian Privacy Principles (APP) [APP]. In FIPP, the most fundamental principle is “notice” [28]. Consumers should be provided with a clear understanding of an entity’s information practices before any personal data is collected from them (e.g., [GDPR Art. 12, GDPR Rectical 58]). In the absence of transparent and understandable notice, it becomes impossible for a consumer to make an informed decision regarding the extent of personal information they choose to disclose.
@@ 2.1 Related Work
Research efforts to improve the presentation and readability of privacy policies have been manifold and encompass various approaches. Kelly et al. [62, 63, 64] introduced an innovative method called privacy nutrition labels (PNLs), designed to facilitate consumers’ understanding of how their information is collected and utilized in a structured table-like way. By employing this concise and organized framework, PNLs effectively aid in understanding the privacy practices of software applications [42, 45, 74, 75, 86]. In a related vein, Cranor [44] proposed a protocol, i.e., Platform for Privacy Preferences, that allows website organizations to declare their intended use of information collected from website users. Alongside these approaches, there has been a growing focus on visual aids such as privacy icons [58, 59, 51]. These small graphical symbols, symbolizing privacy concepts or practices, serve as visual shorthand, simplifying complex privacy policies. By standardizing these icons across various services and platforms, researchers aim to reduce confusion, thereby making privacy policies more readily accessible to the general populace [59].
@@ 2. What data practices are of concern and desire for users?
Administration. The administration often functions as the key regulator within the broader ecosystem of data privacy. A report by the United Nations Conference on Trade and Development reveals that 137 of 194 countries and regions have instituted data protection and privacy legislation [14]. To gain insight into the administrative perspective on these matters, we examine five representative privacy regulations (details are presented in the Appendix 8.1), including the GDPR [8], CCPA [4], the California Online Privacy Protection Act (CalOPPA) [6], the Children’s Online Privacy Protection Act (COPPA) [5], and APP [APP]. These regulations serve as critical touchpoints in our exploration of the administration’s stance on data practices.
@@ 4.1 Context Detection Module
Textual GUI components. To detect and localize privacy-related texts from a given screenshot, we apply a state-of-the-art ultra-lightweight OCR system, named PaddlePaddle-OCR (PP-OCR) [70]. After obtaining these texts, we employ a pre-trained large language model, i.e., GPT-3.5222https://platform.openai.com/docs/models/gpt-3-5, to classify their data types. The specific process is detailed in Appendix LABEL:sec_appendix_gpt. If the answer contains a data type, we will regard the {detected text} as relevant to this data type; otherwise, it does not belong to any data type.
@@ 5.1.1 Context Detection Module
Iconic GUI components. Table 5 shows the comparison of the icon localization capability between our method and previous methods. The results show that our proposed method performs better in identifying the icons curated on GUI screenshots. The failure case analysis is included in the Appendix LABEL:appendix_failure. Table 4b shows the class-wise breakdown for Iconic GUI components, achieving an accuracy of 0.92, a precision of 0.95, and a recall of 0.96. We then compare our icon classification models with several baselines on the RICO-icon dataset. Table 6 shows that our ViT-based model achieves the best performance, but has the largest model size of 328 MB and the slowest time efficiency of 11.28 seconds. The ResNet-based model provides a good balance of performance and efficiency, with the fast time efficiency, decent performance metrics, and a much smaller model size compared to others.
@@ 6.4 Notice-and-Choice Privacy Framework for Mobile Apps: Is This the Way?
While SeePrivacy is not “panacea” for these challenges, it enlightens “a new hope” in this landscape. It suggests a paradigm shift in the privacy notice (what users are agreeing to) and choice (to agree or not) from a generic application level to specific functions or certain GUI components. It is critical to proactively elevate privacy considerations from non-functional requirements to functional requirements in software design and development. Our framework aims to actively prompt users with timely and contextually relevant privacy information, enabling the potential for more granular control in future.
@@ 7 Conclusion and Future Work
Privacy policies have long been criticized for their complexity and poor readability, leading to a growing interest in contextual privacy policies (CPPs) as user-focused alternatives. In this paper, we introduce the concept of CPPs specifically within the domain of mobile applications and propose a novel framework, SeePrivacy, designed to automatically generate these privacy policies. We rigorously evaluated SeePrivacy on our benchmark dataset, Cpp4App, which consists of over 1,200 privacy-related contexts and corresponding privacy policy segments. In the quantitative evaluation, our framework demonstrates 0.88 precision and 0.90 recall to detect contexts; as well as 0.98 precision and 0.96 recall to extract corresponding policy segments. Furthermore, a human evaluation was conducted to validate the functionality our approach, reflected by a median score of 4 on a 5-point Likert scale. Our pre- and post-survey revealed that people exhibit much greater willingness to engage with CPPs, scoring 4.1 out of 5, compared to a mere 2 out of 5 for traditional privacy policies. We also discussed the potential applications in real scenarios, and implications of CPPs and SeePrivacy to the wider community.
Our promising results identify two potential research directions. First, while our current efforts emphasize methodological development and evaluation, we recognize the importance of exploring real-world deployment scenarios. The user experience, particularly how users interact with contextual privacy policies, is vital and could significantly impact the system’s effectiveness. Future work may include integrating SeePrivacy into specific human-device interactions, such as live applications on smartphones or sandbox environments within which applications operate. Second, although we aim to deliver the original privacy policy content without any modification, we notice that some retrieved policy segments are too long to comprehend in a timely manner. Thereby, it is worthwhile to evaluate how to digest lengthy policy segments with minimum information loss and distortion to maintain their integrity and practicability simultaneously.
Last, we believe this work provides a substantial step forward in delivering transparent privacy notices, offering insights into the transition from privacy notices to informed user choices, and paving the way for future research and practical implementation in this crucial area.
@@ Acknowledgments
This research / project is supported by CSIRO’s Data61 PhD Scholarship, and the National Research Foundation, Singapore, and Cyber Security Agency of Singapore under its National Cybersecurity Research and Development Programme, NCRP25-P03-NCR-TAU. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of CSIRO’s Data61, National Research Foundation, Singapore and Cyber Security Agency of Singapore.

### 10 Mar 2024 v2

@@ 1 Introduction
Privacy policies have become the most prevalent privacy notice approach for mobile applications, with the aim of protecting users’ privacy and digital security [53, 87, 56, 64, 30, 55, 37, 70]. On the contrary, application users often find themselves in a privacy paradox, specifically, Digital Resignation. This term refers to a state where users desire to understand and control the information that digital entities own about them, but feel powerless to do so [49]. A study has found that the average length of privacy policies is around 4,000 words for 75 prominent mobile applications and websites, requiring an estimated 16 minutes to read for an average adult [34]. Consequently, according to a survey conducted by Obar and Oeldorf-Hirsch [82], a significant portion of users, approximately 74%, opt for the “quick join” clickwrap option without engaging with privacy policies, primarily due to the lengthy and detailed nature of these documents. These numbers indicate that privacy policies have become increasingly difficult to read and understand in a timely manner, exacerbating the prevalence of digital resignation among users.
@@
In summary, our proposed framework extends the automated generation of CPPs to mobile applications, which are the digital platforms users spend the majority of time on. The key contributions are:
@@ 2 Background
Mobile applications provide users with unparalleled convenience, multiple entertainment options, and various services. At the same time, they are responsible for the collection and exchange of substantial amounts of personal information, raising concerns surrounding data privacy. In response to the potential risks of privacy violations, the U.S. Department of Health, Education, and Welfare proposed a set of Fair Information Practice Principles (FIPP) in 1973 [9, 67]. Later, the Federal Trade Commission adopted these principles as “the notice and choice privacy framework” [82], which is commonly regarded as the foundation for modern privacy regulations such as the European General Data Protection Regulation (GDPR) [10], the California Consumer Privacy Act (CCPA) [6], and the Australian Privacy Principles (APP) [3]. In FIPP, the most fundamental principle is “notice” [27]. Consumers should be provided with a clear understanding of an entity’s information practices before any personal data is collected from them (e.g., [GDPR Art. 12, GDPR Rectical 58]). In the absence of transparent and understandable notice, it becomes impossible for a consumer to make an informed decision regarding the extent of personal information they choose to disclose.
@@ 2.1 Related Work
Research efforts to improve the presentation and readability of privacy policies have been manifold and encompass various approaches. Kelly et al. [61, 62, 63] introduced an innovative method called privacy nutrition labels (PNLs), designed to facilitate consumers’ understanding of how their information is collected and utilized in a structured table-like way. By employing this concise and organized framework, PNLs effectively aid in understanding the privacy practices of software applications [44, 72, 74, 85]. In a related vein, Cranor [43] proposed a protocol, i.e., Platform for Privacy Preferences, that allows website organizations to declare their intended use of information collected from website users. Alongside these approaches, there has been a growing focus on visual aids such as privacy icons [57, 58, 50]. These small graphical symbols, symbolizing privacy concepts or practices, serve as visual shorthand, simplifying complex privacy policies. By standardizing these icons across various services and platforms, researchers aim to reduce confusion, thereby making privacy policies more readily accessible to the general populace [58].
@@ 2. What data practices are of concern and desire for users?
Administration. The administration often functions as the key regulator within the broader ecosystem of data privacy. A report by the United Nations Conference on Trade and Development reveals that 137 of 194 countries and regions have instituted data protection and privacy legislation [16]. To gain insight into the administrative perspective on these matters, we examine five representative privacy regulations, including the GDPR [10], CCPA [6], the California Online Privacy Protection Act (CalOPPA) [8], the Children’s Online Privacy Protection Act (COPPA) [7], and APP [3]. These regulations serve as critical touchpoints in our exploration of the administration’s stance on data practices.
@@ 4.1 Context Detection Module
Textual GUI components. To detect and localize privacy-related texts from a given screenshot, we apply a state-of-the-art ultra-lightweight OCR system, named PaddlePaddle-OCR (PP-OCR) [69]. After obtaining these texts, we employ a pre-trained large language model, i.e., GPT-3.5222https://platform.openai.com/docs/models/gpt-3-5, to classify their data types. If the answer contains a data type, we will regard the {detected text} as relevant to this data type; otherwise, it does not belong to any data type.
@@ 5.1.1 Context Detection Module
Iconic GUI components. Table 5 shows the comparison of the icon localization capability between our method and previous methods. The results show that our proposed method performs better in identifying the icons curated on GUI screenshots. Table 4b shows the class-wise breakdown for Iconic GUI components, achieving an accuracy of 0.92, a precision of 0.95, and a recall of 0.96. We then compare our icon classification models with several baselines on the RICO-icon dataset. Table 6 shows that our ViT-based model achieves the best performance, but has the largest model size of 328 MB and the slowest time efficiency of 11.28 seconds. The ResNet-based model provides a good balance of performance and efficiency, with the fast time efficiency, decent performance metrics, and a much smaller model size compared to others.
@@ 6.4 Notice-and-Choice Privacy Framework for Mobile Apps: Is This the Way?
While SeePrivacy is not a “panacea” for these challenges, it enlightens “a new hope” in this landscape. It suggests a paradigm shift in the privacy notice (what users are agreeing to) and choice (to agree or not) from a generic application level to specific functions or certain GUI components. It is critical to proactively elevate privacy considerations from non-functional to functional requirements in software design and development. Our framework aims to actively prompt users with timely and contextually relevant privacy information, enabling the potential for more granular control in future. In addition, exploring the practice of Right to be Forgotten can sufficiently supplement this approach, empowering users with more control over their choices [106, 107].
@@ 7 Conclusion and Future Work
Privacy policies have long been criticized for their complexity and poor readability, leading to a growing interest in contextual privacy policies (CPPs) as user-focused alternatives. In this paper, we introduce the concept of CPPs specifically within the domain of mobile applications and propose a novel framework, SeePrivacy, designed to automatically generate these privacy policies. We rigorously evaluated SeePrivacy on our benchmark dataset, Cpp4App, which consists of over 1,200 privacy-related contexts and corresponding privacy policy segments. In the quantitative evaluation, our framework demonstrates 0.88 precision and 0.90 recall to detect contexts; as well as 0.98 precision and 0.96 recall to extract corresponding policy segments. Furthermore, a human evaluation was conducted to validate the functionality our approach, reflected by a median score of 4 on a 5-point Likert scale. Our pre- and post-survey revealed that people exhibit much greater willingness to engage with CPPs, scoring 4.1 out of 5, compared to a mere 2 out of 5 for traditional privacy policies. We also discussed the potential adoption scenarios and implications of SeePrivacy to the wider community.
Our promising results identify two potential research directions. First, while our current efforts emphasize methodological development and evaluation, we recognize the importance of exploring real-world deployment scenarios. The user experience, particularly how users interact with contextual privacy policies, is vital and could significantly impact the system’s effectiveness. Future work may include integrating SeePrivacy into specific human-device interactions, such as live applications on smartphones or sandbox environments within which applications operate. Second, although we aim to deliver the original privacy policy content without any modification, we notice that some retrieved policy segments are too long to comprehend in a timely manner. Thereby, it is worthwhile to evaluate how to digest lengthy policy segments with minimum information loss and distortion to maintain their integrity and practicability simultaneously. Overall, we believe this work provides a substantial step forward in delivering transparent privacy notices, paving the way for future research and practical implementation in this crucial area.
@@ Acknowledgments
This research / project is supported by CSIRO’s Data61 PhD Scholarship, and the National Research Foundation, Singapore, and Cyber Security Agency of Singapore under its National Cybersecurity Research and Development Programme, NCRP25-P03-NCR-TAU. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of CSIRO’s Data61, National Research Foundation, Singapore and Cyber Security Agency of Singapore.